Solve the system of equations using substitution.
y = 4x + 10
y = 3x + 7
Please Explain how you go the answer! Thank You!!

Answer by nerdybill(7384)

y = 4x + 10 (equation 1)
y = 3x + 7 (equation 2)
.
Plug definition of y from equation 2 into equation 1 and solve for x:
3x + 7 = 4x + 10
7 = x + 10
-3 = x
.
Plug definition of x above into equation 2 to find y:
y = 3x + 7
y = 3(-3) + 7
y = -9 + 7
y = -2
.
Answer:
x = -3
y = -2